OpenStudy (anonymous):

Double check some of my work please? 1. 2(3) - 3(-2) = 0? 2. The piggy bank contained $25, and $1.50 is added each day. Would the expression of this be written as 25 + (1.50x)? 3. In the first set, the volleyball team made only 8 shots worth one point each. Would the expression for this just be 1 times 8?

OpenStudy (amistre64):

what do 2 negatives equal?

OpenStudy (anonymous):

a positive

OpenStudy (amistre64):

correct, so #1 is a little off

OpenStudy (amistre64):

the others look fine to me

OpenStudy (anonymous):

okay, so the answer is 1??

OpenStudy (amistre64):

you have 3 questions posted, which one are you refering to?

OpenStudy (anonymous):

the first one

OpenStudy (amistre64):

the answer is not 1 then

OpenStudy (anonymous):

oh wait 10, sorry i looked at it wrong

OpenStudy (amistre64):

2(3) = ? -3(-2) = ?? ------ add em up

OpenStudy (anonymous):

oh if the symbol before the 3 changes, the answer is 20?

OpenStudy (amistre64):

3+2 = 5 3 * 2 = 6 :) i think you just did a little err on that lets try 12 altogether ;)

OpenStudy (anonymous):

okay 11?? Sorry this is being so difficult haha

OpenStudy (amistre64):

2(3) - 3(-2) = ? 6 - (-6) = ? 6 + 6 = ?

OpenStudy (anonymous):

12..

OpenStudy (amistre64):

there we go

OpenStudy (anonymous):

Okay, sorry for taking up so much time. Thank you!

OpenStudy (amistre64):

good luck